# PostEdit: Posterior Sampling for Efficient Zero-Shot Image Editing

### v1

@@ A.1 Classifier free diffusion guidance
A.2 Algorithm for Image Reconstruction
@@
  Require: Diffusion model $\boldsymbol{\epsilon}_{\theta}$, step size $h$, image $x_{0}$, measurement $y$, weight $w$, initial prompt $c_{ini}$, encoder ${\mathcal{E}}$, decoder $\mathcal{D}$, optimization steps $N_{L}$.
  $\boldsymbol{z}_{0}\sim\mathcal{E}(\boldsymbol{x}_{0})$, $\boldsymbol{z}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$.
@@
        $\boldsymbol{z}_{0}=\boldsymbol{\hat{z}_{0}}(\boldsymbol{z}_{j},j,c_{ini})$
        $\boldsymbol{z}_{0}^{(j+1)}=(1-w)\cdot\boldsymbol{z}_{0}^{(j)}+w\cdot\boldsymbol{z}_{in}-h\cdot\left(\frac{\|z_{0}^{(j)}-z_{0}\|^{2}}{2\sigma_{t}^{2}}+\frac{\|\mathcal{A}\left(\boldsymbol{z}_{0}^{(j)}\right)-\boldsymbol{y}\|^{2}}{2m^{2}}\right)+\sqrt{2h}\epsilon$,   $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$
@@
     Sample $\boldsymbol{z}_{\tau_{i-1}}\sim\mathcal{N}(\boldsymbol{z}_{0}^{(N_{L})},\sigma_{{\tau_{i-1}}}\boldsymbol{I})$.
@@
The main difference between the reconstruction and editing tasks is the prompt applied to estimate $\boldsymbol{\hat{z}}_{0}$, where it is the initial and target prompts remain the same with each other for image reconstruction.
A.3 Proof of Proposition 2
@@
A.4 Implementation Details
GPU. All our experiments are performed on a single NVIDIA A100 GPU with 80GB memory.
Parameters in Alg. 1. $N$ is set to $5$ for schedule $\{\tau_{i}\}_{i=1}^{N}$. For better quality of editing, we sometimes sample $\boldsymbol{z}_{N}$ by
where $t$ is set to $501$ generally according to the DDPM schedule Ho et al. (2020). The reconstruction process is not restricted to this setting. For Eq. 16, $m$ is set as $0.01$ for both the reconstruction and editing task while $\sigma_{t}$ can be referred to the corresponding timestep of DDPM scheduler Ho et al. (2020). Generally, we apply Eq. 8 for $1$ step to estimate $\boldsymbol{z}_{0}$, and then according to the following schedule to make $\boldsymbol{z}_{\tau_{i}}$ progressively converge to $\boldsymbol{z}_{0}$.
$\{\tau_{i}\}_{i=1}^{5}=\{501,401,301,201,101,1\}.$ (24)
The parameter $w$ is usually set to a minimal value such as $0.05$ for most cases or 00 and $0.1$ for easy and hard cases. Additionally, $h$ is always set to 1e-5 for image edting and reconstruction tasks.
Parameters of Consistency models. $k_{skip}$ and $k_{out}$ shown in Eq. 8 are set to 0 and 1 for most cases respectively.
ODE Solvers. We adopt the solver of LCM Luo et al. (2023a) distilled from Dreamshaper v7 fine-tune of Stable-Diffusion v1-5 for images editing task. For reconstruction, different solvers, for instance, DDIM Song et al. (2021a), DDPM Ho et al. (2020), and Song et al. (2023) based on Stable Diffusion Rombach et al. (2022) are able to work out satisfied reconstruction quality.
Oversampling Matrix. We set $\sigma$ shown in Eq. 27 to $0.01$ and use an oversampling factor $k=2$ and $n=8$.
2D Discrete Fourier Transform Matrix. The 2D Fourier transform is defined as
$u=0,1,...,M-1;\ \ v=0,1,...,N-1,$
Probability of Masked Features. We use the probability equal to 0.5 for a randomly mask process, which represents whether one of the latent features is masked or not.
Measurement $\boldsymbol{y}$ Used for Image Reconstruction. The measurement $\boldsymbol{y}$ can be chosen from any of the following two, which are defined as linear and nonlinear operations relating to initial image $\boldsymbol{z}_{0}$ in latent space
where $\boldsymbol{F}$ and $\boldsymbol{P}$ denote the 2D discrete Fourier transform matrix and the oversampling matrix with ratio $k/n$ respectively for Eq. 27.
SD Model. We adopt LCM-SD1.5 for all the experiments Luo et al. (2023a).
FFHQ Model. We adopt the ffhq_10m.pt with a size of 357.1MB as the baseline model for all the experiments relating to the FFHQ dataset.
A.5 Comparison The Images Layout of Different Datasets
Figure 6: Layouts Comparison of Evaluated Output for the same Objects in Different Intermediate Timesteps.
Figure 7: Layouts of Evaluated Output for Different Objects and Timesteps.
A.6 Comparison Between LCM-SD1.5 And SDXL-Turbo
@@ A.5 Comparison The Images Layout of Different Datasets
A.7 Additional Results
The additional experimental results are exhibited here.
Figure 9: Additional Comparison Results.
Figure 10: Additional reconstruction Results.

### v2

@@ A.14 Classifier free diffusion guidance
A.15 Algorithm for Image Reconstruction
@@
  Require: Diffusion model $\boldsymbol{\epsilon}_{\theta}$, diffusion sampler $\boldsymbol{\hat{z}_{0}}\left(\cdot\right)$, posterior sampling steps $N$, step size $h$, image $x_{0}$, measurement $y$, weight $w$, initial prompt $c_{ini}$, encoder ${\mathcal{E}}$, decoder $\mathcal{D}$, noise schedule $\alpha(t)$, $\sigma(t)$, optimization steps $N_{L}$ and posterior sampler sequence $\{\tau_{i}\}_{i=0}^{N}$.
  $\boldsymbol{z}_{\tau_{N}}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$.
@@
        $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$.
        $\boldsymbol{z}_{0}^{(j+1)}=(1-w)\cdot\boldsymbol{z}_{0}^{(j)}+w\cdot\boldsymbol{z}_{in}-h\cdot\nabla_{\boldsymbol{z}_{0}^{(k)}}\left(\frac{\|z_{0}^{(j)}-z_{0}\|^{2}}{2\sigma_{t}^{2}}+\frac{\|\mathcal{A}\left(\boldsymbol{z}_{0}^{(j)}\right)-\boldsymbol{y}\|^{2}}{2m^{2}}\right)+\sqrt{2h}\boldsymbol{\epsilon}$.
@@
     Sample $\boldsymbol{z}_{\tau_{i-1}}\sim\mathcal{N}\left(\boldsymbol{z}_{0}^{(N_{L})},\sigma_{{\tau_{i-1}}}\boldsymbol{I}\right)$.
@@
The differences between the reconstruction and editing tasks is the ODE solvers applied to estimate $\boldsymbol{\hat{z}}_{0}$, and the initial and target prompts remain the same with each other for image reconstruction. All the parameter Settings are shown specifically in App. A.2.
A.16 Proof of Proposition 2
@@
Figure 6: Layouts of evaluated outputs for the same objects at different intermediate timesteps.
Figure 7: Layouts of evaluated output for various objects and timesteps.
Figure 8: Example of Long-text Editing.
Figure 9: Example of Long-text Editing.
Figure 10: Example of Long-text Editing.
Figure 11: Example of Long-text Editing.
Figure 12: Additional Comparison Results of The Remained Baselines.
Figure 13: Intermediate Results.
Figure 14: Additional Comparison Results Based on Different Measurements.
Figure 15: Additional reconstruction Results.
Figure 16: Additional reconstruction Results.
@@
Figure 18: Instruction Generation process GPT.
Figure 19: Comparison of Different methods for Reconstruction of High Frequency Details. $\dagger$ represents 10000 optimization steps are adopted for this result.